OpenStudy (anonymous):

Which is the exact value of the expression the square root of 48. − the square root of 75. + the square root of 192.

OpenStudy (anonymous):

Just need help.

OpenStudy (anonymous):

\[\sqrt{48} - \sqrt{75} + \sqrt{192}\]

OpenStudy (anonymous):

\[7\sqrt{3}\]

OpenStudy (anonymous):

@Data_LG2 I need to learn how. I need steps. I need to understand it.

OpenStudy (anonymous):

@evilmath Thank you (:

OpenStudy (anonymous):

\[\sqrt{16 \times 3 } + \sqrt{\frac{ 25 }{ 2 } \times 6} + \sqrt{64 \times 3 }\]

OpenStudy (anonymous):

do you follow?

OpenStudy (anonymous):

Woops I meant to write \[4\sqrt{3} - 5\sqrt{3} + 8\sqrt{3}\]